# Event Camera Meets Resource-Aware Mobile Computing:

### v3

@@ 2.4. Challenges of event cameras in mobile sensing
Since event cameras operate fundamentally differently from frame-based cameras by capturing per-pixel brightness changes asynchronously as events, their integration into mobile devices poses several challenges:
@@
(i) How to mitigate event bursts and accurately extract features from event data, given the high-speed operation of mobile devices and event cameras’ lack of stable, persistent semantic information? Event cameras are highly sensitive to illumination changes, with even minor variations triggering numerous events. On high-agility mobile devices, rapid scene changes captured by onboard event cameras can trigger event bursts, generating thousands of events in a short time. Taking the DAVIS 346 event camera (346 × 260 resolution) as an example, let $v$ denote the translational velocity of the drone and $w$ its angular velocity. Under normal flight conditions ($2\text{m/s}\leq v\leq 4\text{m/s}$, $5^{\circ}/\text{s}\leq w\leq 15^{\circ}/\text{s}$), the event generation rate is 298 $e/ms$. During rapid translation ($18\text{m/s}\leq v\leq 25\text{m/s}$, $5^{\circ}/\text{s}\leq w\leq 15^{\circ}/\text{s}$), the rate increases to 945 $e/ms$, while in rapid rotation ($2\text{m/s}\leq v\leq 4\text{m/s}$, $75^{\circ}/\text{s}\leq w\leq 95^{\circ}/\text{s}$), it further rises to 1437 $e/ms$ [39]. In contrast, the data rate of a frame-based camera is independent of motion dynamics and scene content. Regardless of whether the scene is static or highly dynamic, frames are sampled at a fixed rate, e.g., 1MP × 30fps = 30M pixels/s, resulting in a stable and uniform data stream. This comparison highlights the non-uniform nature of event data: static scenes generate almost no events, while high-speed motion can trigger explosive outputs far exceeding frame rates. This sparsity–burst duality, coupled with the lack of stable texture or semantic cues, makes efficient processing difficult, as meaningful signals are often buried in motion-induced noise.
@@
(ii) How to efficiently process a large volume of event data given on-board constrained resources? Mobile devices typically rely on low-power embedded systems for efficiency, inherently limiting their computational capacity. For example, a smartphone SoC such as the Qualcomm Snapdragon 8Gen3 has a typical power budget of 5–7WW, with its integrated NPU delivering around 30TOPSTOPS at INT8 precision. By contrast, a laptop equipped with an Intel Core i7-13700H can consume up to 45WW per CPU package and provide 1–2 $TFLOPS$ of FP32 compute power. For high-performance drones, the NVIDIA Jetson Orin NX module consumes 15–25WW and delivers 100 $TOPS$ of AI compute, optimized for onboard vision tasks. In comparison, a NVIDIA GeForce RTX 4060Ti (160WW) offers up to 22 $TFLOPS$ FP32 and 288 $TOPS$ INT8, while the RTX 4090 (450WW) reaches 83 $TFLOPS$ FP32 and over 1.3 $POPS$ INT8. Clearly, mobile devices trade computational resources for energy efficiency. Meanwhile, onboard event cameras capture rapid scene changes, generating large volumes of events that demand efficient processing. For instance, the iniVation DVXplorer can reach peak output rates above 1MepsMeps, the Sony IMX636 sensor used in Prophesee Gen4.1 exceeds 10MepsMeps, and the high-speed CelePixel Taurus supports up to 240MepsMeps. Such workloads heavily strain limited computational resources. In a typical scenario, an algorithm processing a 5MepsMeps event stream with 100 $GOP/s$ demand is trivial for an RTX 4060Ti (288 $TOPS$), leaving ample headroom for other tasks. For a smartphone NPU (30 $TOPS$), however, achieving low latency and high energy efficiency is challenging due to memory sharing constraints and thermal throttling.
2.5. Commercial product & comparation
@@
Event cameras are increasingly entering the commercial market, with several companies offering distinct products. Representative suppliers include iniVation, Prophesee, Lucid Vision Labs and CelePixel. Their commercial offerings emphasize features relevant to mobile and embedded contexts, such as low-power operation, compact modules, mobile-friendly interfaces, or robust industrial connectivity.
@@
Table 3. Specifications of Event Camera Models for Mobile/Embedded Sensing Applications
@@
iniVation DVXplorer 640×480 90-120 9.0 Max 12W Multi-camera synchronization; Robust aluminum casing
DVXplorer Micro 640×480 110 9.0 <140mA @ 5VDC Compact and lightweight design
@@
DAVIS346 346×260 120 18.5 Typical 180mA @ 5VDC Hybrid output (provides both event and frame data)
@@
DAVIS346 AER 346×260 120 18.5 Typical 180mA @ 5VDC Integrated IMU for self-contained visual-inertial sensing
@@
Prophesee EVK4 HD 1280×720 >86 4.86×4.86 Typical 0.5W High-resolution sensor with external trigger support
EVK5 HD 1280×720 >110 4.86×4.86 Typical 0.5W On-board advanced processing; Hardware trigger support
@@
GENX320 320×320 >120 6.3 3mW Ultra-low power consumption; Compact, mobile-optimized design
@@
CelePixel CeleX-V 1280×800 - 9.8 400mW Multi-mode output (Event, Grayscale, Accumulated frames)
@@
CeleX5-MIPI 1280×800 - 9.8 - Designed for direct, low-level System-on-Chip (SoC) integration
@@
Lucid Vision Labs TRT009S-EC 1280×720 120 4.86 - Industrial-grade robustness and reliable data streaming
@@
Inivation is a prominent leader in the event camera industry, specializes in developing high-resolution, energy-efficient event cameras [47]. Inivation’s flagship sensors, like the DAVIS240 and DAVIS346, combine event-based and frame-based data, enabling real-time analysis of both modalities. With resolutions up to 1 megapixel, they rank among the highest-resolution event cameras available. Designed for excellent low-light performance and outdoor durability, these sensors are ideal for applications in robotics, autonomous systems, and more.
@@
Prophesee is a leading innovator in event-based vision technology, renowned for its cameras’ ultra-high dynamic range exceeding 120 dB [48]. Their flagship Metavision sensors, including the EVK4 HD and EVK5 HD cameras, offer microsecond-level temporal resolution for applications requiring rapid response and high accuracy, such as autonomous driving and industrial automation. Recently, they have introduced the GENX320, a new sensor designed with features particularly beneficial for mobile sensing applications. With its compact 320×320 resolution, small 6.3µm pixels, and ultra-low power consumption, the GENX320 addresses key constraints in battery-powered mobile devices, making it well-suited for applications such as AR/VR headsets, drones, and other embedded platforms [46].
@@
Lucid Vision Labs, a manufacturer of industrial cameras [49], offers the TRT009S-EC Triton camera built on the Sony IMX636 event sensor. It utilizes a GigE Vision interface with Power over Ethernet (PoE) to ensure robust and reliable data streaming. This design prioritizes operational stability for mobile robotic platforms over the ultra-low power consumption typical of other mobile-optimized sensors.
@@
CelePixel developed smart sensory platforms with a unique on-chip processing architecture. While its official website is inactive, technical resources remain available via software repositories. Their CeleX-V sensor [44] is notable for a multi-mode capability, outputting events, full frames, and hardware-generated event-accumulation views. The CeleX5-MIPI [45] variant targets mobile integration, providing a MIPI CSI-2 interface for low-power connection.
@@
Comparation of products. Table 3 compares commercial event cameras, revealing significant diversity in their specifications. Resolutions range from 320x320 to high-definition 1280x800 pixels, while a high dynamic range (often >110 dB) is a common strength for challenging lighting. A critical differentiator is power consumption, which spans from just 3 mW for mobile-optimized models to 12 W for high-performance ones. This reflects divergent design goals, from ultra-low-power, compact devices to performance-focused industrial models with advanced features like integrated IMUs, hybrid outputs, and multi-camera synchronization for complex robotic systems.
@@ 2.6. Event-based datasets
DVS-Pedestrian [51]. The DVS-Pedestrian dataset is a benchmark dataset designed specifically for pedestrian detection and tracking, captured using a DVS in various urban environments. It features sequences of pedestrians performing various actions (walking, running, standing) under different lighting conditions and backgrounds.
@@
VECtor [55]. The VECtor Event Dataset is the first SLAM benchmark dataset captured using a fully synchronized multi-sensor setup, including event-based and regular stereo cameras, RGB-D sensors, LiDAR, and IMU, with complete 6-DoF ground truth for diverse scenarios. It captures the full spectrum of motion dynamics and environmental conditions while providing precise calibration, specifically designed to address challenges unique to dynamic vision sensors.
@@
DVS-Pedestrian [51] 2019 0.1 hours, 4.6K annotations Dynamic Pedestrians Daytime 4.6K Walking street
@@
1 Mpx Automotive [56] 2020 15 hours Dynamic Cars, pedestrians, two-wheelers Daytime, night 25M bboxes Object Detection
@@
FE108 [57] 2021 1.5 hours Dynamic 21 object types LL, HDR, fast motion 208K frames Object Tracking
VECtor [55] 2022 12 video clips Dynamic Static objects Standard light, low-light, HDR - SLAM
eTraM [58] 2024 10 hours, 2M annotations Static Vehicles, pedestrians, micro-mobility Daytime, night, twilight 2M Intersections, roadways, streets
LLE-VOS[59] 2024 70 video clips Dynamic Pedestrians, other targets Normal, low-light 5600 Gym, classroom, zoo
REVD [60] 2024 21 sequences Dynamic Various scenes Varied 21 paired seq. Video Deblurring
SDE [61] 2024 91 sequences Dynamic Indoor, outdoor scenes Low-light, normal 30K+ pairs Low-Light Enhancement
eTraM [58]. The eTraM (Event-based Traffic Monitoring) dataset a fully event-based traffic perception benchmark captured using a high-resolution Prophesee EVK4 HD event camera. It features annotated traffic data with diverse vehicles, pedestrians, and micro-mobility objects under challenging lighting and weather conditions, including high glare, overexposure, underexposure, nighttime, twilight, and rainy days.
LLE-VOS [59]. The LLE-VOS (Low-Light Event-based Video Object Segmentation) dataset is designed for event-based video object segmentation in low-light conditions, providing synchronized event and frame data with ground-truth masks. It captures challenging scenarios including night-time and indoor scenes.
FE108 [57]. The FE108 dataset is a large-scale, frame-event-based resource for object tracking. It comprises 108 sequences with 21 object types under challenging conditions like low light, HDR, and fast motion. With high-frequency ground truth for both domains, it is highly suitable for evaluating multi-modal tracking algorithms.
1 Mpx Automotive Detection Dataset [56]. This dataset is the first large-scale, high-resolution benchmark for automotive object detection. It provides over 14 hours of recordings with more than 25 million bounding box annotations for cars, pedestrians, and two-wheelers, making it ideal for training robust detectors for autonomous driving.
REVD [60]. The Real-world Event Video Deblurring (REVD) dataset is the first real-world benchmark for its task. It offers synchronized high-resolution blurred videos, sharp ground-truth counterparts, and event streams captured in scenes with extreme motion blur, serving as a critical resource for video restoration algorithms.
SDE dataset [61]. The SDE dataset is a large-scale, real-world benchmark for low-light image enhancement. It consists of over 30,000 spatially and temporally aligned image-event pairs captured in varied lighting conditions. This precise alignment enables the development of robust enhancement techniques for real-world scenarios.
@@ 2.7. Synthetic data generation: simulators and approaches
While deep learning has advanced event-based vision, progress depends on large, diverse datasets, yet collecting dense ground truth for tasks like optical flow is costly [27]. Simulators address this by generating synthetic event data with pixel-accurate annotations in controllable environments, enabling systematic training and evaluation. This section reviews major simulation approaches, categorized as physics-based rendering, video-to-event conversion, neural rendering, and AI-driven generation, as well as quality assessment methods (Tab. 5).
Physics-based rendering and simulation This approach typically involves rendering high-framerate video from a 3D environment and then converting these frames to events based on a pixel model. A pioneering tool, ESIM [62], synthesizes events by interpolating logarithmic intensity changes to provide “perfect” data with microsecond resolution. For large-scale generation, frameworks like InteriorNet [63] create photorealistic indoor scenes with event streams and corresponding multi-modal data (RGB-D, IMU, semantics). In contrast, some simpler methods just threshold frame differences, a faster but less precise technique [64].
Video-to-event conversion To leverage existing video archives, video-to-event (V2E) converters transform standard video into event streams. The v2e toolbox [65], for instance, implements a sophisticated pixel model that accounts for non-ideal behaviors like bandwidth limits, threshold mismatch, and noise, yielding highly realistic DVS events. Other robust tools, including the V2CE Toolbox [66] and the Prophesee Video to Event Simulator [67], also provide robust methods for converting frame-based video into the event domain.
Neural rendering and AI-driven event generation More recently, neural rendering and generative AI have emerged. Neural Radiance Fields (NeRFs) offer a new direction; EvDNeRF [68], the first dynamic NeRF trained on events, can synthesize event streams of dynamic scenes from novel viewpoints. Pushing this further, generative AI models like Text-to-Events [69] bypass intermediate frames entirely, using a latent diffusion model to generate event data directly from text prompts, though currently limited to specific domains like gestures.
Differentiable simulators and quality assessment Recent research thrusts are differentiable simulation and quality assessment. SENPI [70] exemplifies the former, offering a fully differentiable framework for end-to-end co-optimization of sensor parameters and a network’s architecture. To address sim-to-real gap, Event Quality Score (EQS) [71] provides a metric by comparing latent features from a network processing real versus synthetic data. The resulting score correlates with real-world performance and can serve as a loss function to improve simulators.

### v4

@@ 2.4. Challenges of event cameras in mobile embodied perception
Since event cameras operate fundamentally differently from frame-based cameras by capturing per-pixel brightness changes asynchronously as events, their integration into mobile agents poses several challenges:
@@
(i) How to mitigate event bursts and accurately extract features from event data, given the high-speed operation of mobile agents and event cameras’ lack of stable, persistent semantic information? Event cameras are highly sensitive to illumination changes, with even minor variations triggering numerous events. On high-agility mobile agents, rapid scene changes captured by onboard event cameras can trigger event bursts, generating thousands of events in a short time. Taking the DAVIS 346 event camera (346 × 260 resolution) as an example, let $v$ denote the translational velocity of the drone and $w$ its angular velocity. Under normal flight conditions ($2\text{m/s}\leq v\leq 4\text{m/s}$, $5^{\circ}/\text{s}\leq w\leq 15^{\circ}/\text{s}$), the event generation rate is 298 $e/ms$. During rapid translation ($18\text{m/s}\leq v\leq 25\text{m/s}$, $5^{\circ}/\text{s}\leq w\leq 15^{\circ}/\text{s}$), the rate increases to 945 $e/ms$, while in rapid rotation ($2\text{m/s}\leq v\leq 4\text{m/s}$, $75^{\circ}/\text{s}\leq w\leq 95^{\circ}/\text{s}$), it further rises to 1437 $e/ms$ [xu2023taming]. In contrast, the data rate of a frame-based camera is independent of motion dynamics and scene content. Regardless of whether the scene is static or highly dynamic, frames are sampled at a fixed rate, e.g., 1MP × 30fps = 30M pixels/s, resulting in a stable and uniform data stream. This comparison highlights the non-uniform nature of event data: static scenes generate almost no events, while high-speed motion can trigger explosive outputs far exceeding frame rates. This sparsity–burst duality, coupled with the lack of stable texture or semantic cues, makes efficient processing difficult, as meaningful signals are often buried in motion-induced noise.
@@
(ii) How to efficiently process a large volume of event data given on-board constrained resources? Mobile agents typically rely on low-power embedded systems for efficiency, inherently limiting their computational capacity. For example, a smartphone SoC such as the Qualcomm Snapdragon 8Gen3 has a typical power budget of 5–7WW, with its integrated NPU delivering around 30TOPSTOPS at INT8 precision. By contrast, a laptop equipped with an Intel Core i7-13700H can consume up to 45WW per CPU package and provide 1–2 $TFLOPS$ of FP32 compute power. For high-performance drones, the NVIDIA Jetson Orin NX module consumes 15–25WW and delivers 100 $TOPS$ of AI compute, optimized for onboard vision tasks. In comparison, a NVIDIA GeForce RTX 4060Ti (160WW) offers up to 22 $TFLOPS$ FP32 and 288 $TOPS$ INT8, while the RTX 4090 (450WW) reaches 83 $TFLOPS$ FP32 and over 1.3 $POPS$ INT8. Clearly, mobile agents trade computational resources for energy efficiency. Meanwhile, onboard event cameras capture rapid scene changes, generating large volumes of events that demand efficient processing. For instance, the iniVation DVXplorer can reach peak output rates above 1MepsMeps, the Sony IMX636 sensor used in Prophesee Gen4.1 exceeds 10MepsMeps, and the high-speed CelePixel Taurus supports up to 240MepsMeps. Such workloads heavily strain limited computational resources. In a typical scenario, an algorithm processing a 5MepsMeps event stream with 100 $GOP/s$ demand is trivial for an RTX 4060Ti (288 $TOPS$), leaving ample headroom for other tasks. For a smartphone NPU (30 $TOPS$), however, achieving low latency and high energy efficiency is challenging due to memory sharing constraints and thermal throttling.
2.5. Commercial product & comparison
@@
Event cameras are increasingly entering the commercial market, with several companies offering distinct products. Representative suppliers include iniVation, Prophesee, Lucid Vision Labs and CelePixel. Their commercial offerings emphasize features relevant to mobile embodied perception contexts, such as low-power operation, compact modules, mobile-friendly interfaces, or robust industrial connectivity.
@@
Table 3. Specifications of event camera models for mobile embodied perception applications
@@
iniVation DVXplorer 640×480 90-120 9.0 Max 12W Multi-camera synchronization; Robust aluminum casing.
DVXplorer Micro 640×480 110 9.0 <140mA @ 5VDC Compact and lightweight design.
@@
DAVIS346 346×260 120 18.5 Typical 180mA @ 5VDC Hybrid output (provides both event and frame data).
@@
DAVIS346 AER 346×260 120 18.5 Typical 180mA @ 5VDC Integrated IMU for self-contained visual-inertial perception.
@@
Prophesee EVK4 HD 1280×720 >86 4.86×4.86 Typical 0.5W High-resolution sensor with external trigger support.
EVK5 HD 1280×720 >110 4.86×4.86 Typical 0.5W On-board advanced processing; Hardware trigger support.
@@
GENX320 320×320 >120 6.3 3mW Ultra-low power consumption; Compact, mobile-optimized design.
@@
CelePixel CeleX-V 1280×800 - 9.8 400mW Multi-mode output (Event, Grayscale, Accumulated frames).
@@
CeleX5-MIPI 1280×800 - 9.8 - Designed for direct, low-level System-on-Chip (SoC) integration.
@@
Lucid Vision Labs TRT009S-EC 1280×720 120 4.86 - Industrial-grade robustness and reliable data streaming.
@@
Inivation is a prominent leader in the event camera industry, specializes in developing high-resolution, energy-efficient event cameras [inivation]. Inivation’s flagship sensors, like the DAVIS240 and DAVIS346, combine event-based and frame-based data, enabling real-time analysis of both modalities. With resolutions up to 1 megapixel, they rank among the highest-resolution event cameras available. Designed for excellent low-light performance and outdoor durability, these sensors are ideal for applications in robotics, autonomous systems, and more.
@@
Prophesee is a leading innovator in event-based vision technology, renowned for its cameras’ ultra-high dynamic range exceeding 120 dB [prophesee]. Their flagship Metavision sensors, including the EVK4 HD and EVK5 HD cameras, offer microsecond-level temporal resolution for applications requiring rapid response and high accuracy, such as autonomous driving and industrial automation. Recently, they have introduced the GENX320, a new sensor designed with features particularly beneficial for mobile embodied perception applications. With its compact 320×320 resolution, small 6.3µm pixels, and ultra-low power consumption, the GENX320 addresses key constraints in battery-powered mobile agents, making it well-suited for applications such as AR/VR headsets, drones, and other embedded agents [prophesee_genx320_datasheet].
@@
Lucid Vision Labs, a manufacturer of industrial cameras [lucid_vision_labs_cn], offers the TRT009S-EC Triton camera built on the Sony IMX636 event sensor. It utilizes a GigE Vision interface with Power over Ethernet (PoE) to ensure robust and reliable data streaming. This design prioritizes operational stability for mobile robotic agents over the ultra-low power consumption typical of other mobile-optimized sensors.
@@
CelePixel developed smart sensory platforms with a unique on-chip processing architecture. While its official website is inactive, technical resources remain available via software repositories. Their CeleX-V sensor [chen2019live] is notable for a multi-mode capability, outputting events, full frames, and hardware-generated event-accumulation views. The CeleX5-MIPI [celex5mipi_datasheet] variant targets mobile integration, providing a MIPI CSI-2 interface for low-power connection.
@@
Comparation of products. Tab. 3 compares commercial event cameras, revealing significant diversity in their specifications. Resolutions range from 320x320 to high-definition 1280x800 pixels, while a high dynamic range (often >110 dB) is a common strength for challenging lighting. A critical differentiator is power consumption, which spans from just 3 mW for mobile-optimized models to 12 W for high-performance ones. This reflects divergent design goals, from ultra-low-power, compact agents to performance-focused industrial models with advanced features like integrated IMUs, hybrid outputs, and multi-camera synchronization for complex robotic systems.
@@ 2.6. Event-based datasets
DVS-Pedestrian [sakhai2024pedestrianintentionpredictionadverse]. The DVS-Pedestrian dataset is a benchmark dataset designed specifically for pedestrian detection and tracking, captured using a DVS in various urban environments. It features sequences of pedestrians performing various actions (walking, running, standing) under different lighting conditions and backgrounds.
@@
VECtor [Gao_2022]. The VECtor Event Dataset is the first SLAM benchmark dataset captured using a fully synchronized multi-sensor setup, including event-based and regular stereo cameras, RGB-D sensors, LiDAR, and IMU, with complete 6-DoF ground truth for diverse scenarios. It captures the full spectrum of motion dynamics and environmental conditions while providing precise calibration, specifically designed to address challenges unique to dynamic vision sensors.
@@
DVS-Pedestrian [sakhai2024pedestrianintentionpredictionadverse] 2019 0.1 hours, 4.6K annotations Dynamic Pedestrians Daytime 4.6K Walking street
@@
1 Mpx Automotive [perot2020learning] 2020 15 hours Dynamic Cars, pedestrians, two-wheelers Daytime, night 25M bboxes Object Detection
@@
FE108 [zhang2021object] 2021 1.5 hours Dynamic 21 object types LL, HDR, fast motion 208K frames Object Tracking
VECtor [Gao_2022] 2022 12 video clips Dynamic Static objects Standard light, low-light, HDR - SLAM
eTraM [verma2024etrameventbasedtrafficmonitoring] 2024 10 hours, 2M annotations Static Vehicles, pedestrians, micro-mobility Daytime, night, twilight 2M Intersections, roadways, streets
LLE-VOS[li2024event] 2024 70 video clips Dynamic Pedestrians, other targets Normal, low-light 5600 Gym, classroom, zoo
REVD [kim2024frequency] 2024 21 sequences Dynamic Various scenes Varied 21 paired seq. Video Deblurring
SDE [liang2024towards] 2024 91 sequences Dynamic Indoor, outdoor scenes Low-light, normal 30K+ pairs Low-Light Enhancement
eTraM [verma2024etrameventbasedtrafficmonitoring]. The eTraM (Event-based Traffic Monitoring) dataset a fully event-based traffic perception benchmark captured using a high-resolution Prophesee EVK4 HD event camera. It features annotated traffic data with diverse vehicles, pedestrians, and micro-mobility objects under challenging lighting and weather conditions, including high glare, overexposure, underexposure, nighttime, twilight, and rainy days.
LLE-VOS [li2024event]. The LLE-VOS (Low-Light Event-based Video Object Segmentation) dataset is designed for event-based video object segmentation in low-light conditions, providing synchronized event and frame data with ground-truth masks. It captures challenging scenarios including night-time and indoor scenes.
FE108 [zhang2021object]. The FE108 dataset is a large-scale, frame-event-based resource for object tracking. It comprises 108 sequences with 21 object types under challenging conditions like low light, HDR, and fast motion. With high-frequency ground truth for both domains, it is highly suitable for evaluating multi-modal tracking algorithms.
1 Mpx Automotive Detection Dataset [perot2020learning]. This dataset is the first large-scale, high-resolution benchmark for automotive object detection. It provides over 14 hours of recordings with more than 25 million bounding box annotations for cars, pedestrians, and two-wheelers, making it ideal for training robust detectors for autonomous driving.
REVD [kim2024frequency]. The Real-world Event Video Deblurring (REVD) dataset is the first real-world benchmark for its task. It offers synchronized high-resolution blurred videos, sharp ground-truth counterparts, and event streams captured in scenes with extreme motion blur, serving as a critical resource for video restoration algorithms.
SDE dataset [liang2024towards]. The SDE dataset is a large-scale, real-world benchmark for low-light image enhancement. It consists of over 30,000 spatially and temporally aligned image-event pairs captured in varied lighting conditions. This precise alignment enables the development of robust enhancement techniques for real-world scenarios.
@@ 2.7. Synthetic data generation: simulators and approaches
While deep learning has advanced event-based vision, progress depends on large, diverse datasets, yet collecting dense ground truth for tasks like optical flow is costly [gallego2020event]. Simulators address this by generating synthetic event data with pixel-accurate annotations in controllable environments, enabling systematic training and evaluation. This section reviews major simulation approaches, categorized as physics-based rendering, video-to-event conversion, neural rendering, and AI-driven generation, as well as quality assessment methods (Tab. 5).
@@
Physics-based rendering and simulation. This approach typically involves rendering high-framerate video from a 3D environment and then converting these frames to events based on a pixel model. A pioneering tool, ESIM [mueggler2017event], synthesizes events by interpolating logarithmic intensity changes to provide “perfect” data with microsecond resolution. For large-scale generation, frameworks like InteriorNet [li2018interiornet] create photorealistic indoor scenes with event streams and corresponding multi-modal data (RGB-D, IMU, semantics). In contrast, some simpler methods just threshold frame differences, a faster but less precise technique [kaiser2016towards].
Video-to-event conversion. To leverage existing video archives, video-to-event (V2E) converters transform standard video into event streams. The v2e toolbox [hu2021v2e], for instance, implements a sophisticated pixel model that accounts for non-ideal behaviors like bandwidth limits, threshold mismatch, and noise, yielding highly realistic DVS events. Other robust tools, including the V2CE Toolbox [zhang2024v2ce] and the Prophesee Video to Event Simulator [propheseemetavision11], also provide robust methods for converting frame-based video into the event domain.
Neural rendering and AI-driven event generation. More recently, neural rendering and generative AI have emerged. Neural Radiance Fields (NeRFs) offer a new direction; EvDNeRF [bhattacharya2024evdnerf], the first dynamic NeRF trained on events, can synthesize event streams of dynamic scenes from novel viewpoints. Pushing this further, generative AI models like Text-to-Events [ott2024text] bypass intermediate frames entirely, using a latent diffusion model to generate event data directly from text prompts, though currently limited to specific domains like gestures.
Differentiable simulators and quality assessment. Recent research thrusts are differentiable simulation and quality assessment. SENPI [greene2025pytorch] exemplifies the former, offering a fully differentiable framework for end-to-end co-optimization of sensor parameters and a network’s architecture. To address sim-to-real gap, Event Quality Score (EQS) [chanda2025event] provides a metric by comparing latent features from a network processing real versus synthetic data. The resulting score correlates with real-world performance and can serve as a loss function to improve simulators.